OpenStudy (anonymous):

Please help I'm struggling..... Captown is the capital city of Maltania. If the population of Captown is 25 percent of the rest of the population of Maltania, then the population of Captown is what percent of the entire population of Maltania?

OpenStudy (anonymous):

You will want to find c/m c=pop of captown m=pop of maltania m-c=rest of population c/m-c=25/100

OpenStudy (anonymous):

why is it m-c?

OpenStudy (anonymous):

@nincompoop

OpenStudy (ribhu):

let captown population be c

OpenStudy (ribhu):

maltania population be m

OpenStudy (anonymous):

whats the difference between rest of the population and the entire population?

OpenStudy (anonymous):

M= Maltania - C= Captown

OpenStudy (ribhu):

remaining population of maltania = m-c

OpenStudy (ribhu):

so c = 0.25(m-c) 1.25c=m

OpenStudy (ribhu):

c/m = (1/1.25)*100

OpenStudy (ribhu):

80

OpenStudy (anonymous):

i dont understand what the difference between rest of the popualation and the entire population is....

OpenStudy (ribhu):

see rest of the population is when population of captown is subtracted from population of maltania

OpenStudy (ribhu):

entire population is population of maltania = m

OpenStudy (anonymous):

ohhh ...

OpenStudy (ribhu):

@yomamabf i hope this will make u clear

OpenStudy (anonymous):

so the entire population is including captown?

OpenStudy (ribhu):

yeah. thats what the calculation shows

OpenStudy (anonymous):

ohhhh kay thanks

OpenStudy (godlovesme):

@yomamabf can u please close this if u r done with the post please? thanks :)